# End-to-end Autonomous Driving: Challenges and Frontiers

### v1

@@ 3 Benchmarking
Autonomous driving systems require a comprehensive evaluation of their reliability to ensure safety [86, 87]. To accomplish this, researchers must benchmark these systems using appropriate datasets, simulators, and metrics. This section delineates the two methods of large-scale benchmarking for end-to-end autonomous driving systems: (1) online or closed-loop evaluation in simulation, and (2) offline or open-loop evaluation on human driving datasets. We focus, in particular, on the more principled online setting and provide a brief summary of offline evaluation for completeness.
3.1 Online Evaluation (Closed-loop)
Conducting tests of self-driving systems in the real world is costly and risky. To address this challenge, simulation is a viable alternative [88, 89, 77, 90, 91, 92, 14]. Simulators facilitate rapid prototyping and testing, enable the quick iteration of ideas, and provide low-cost access to a wide range of scenarios. In addition, simulators offer tools for measuring performance reliably and accurately. However, their primary disadvantage is that the results obtained in a simulated environment do not necessarily generalize to the real world (Sec. 4.8.3).
Closed-loop evaluation involves constructing a simulated environment that closely mimics a real-world driving environment. The evaluation of the driving system entails deploying the system in the simulated environment and measuring its performance over time. The system has to navigate safely through simulated traffic while progressing toward a designated goal location. There are three main sub-tasks involved in developing such simulators for evaluation: parameter initialization, traffic simulation, and sensor simulation. We briefly describe these sub-tasks below, followed by a summary of currently available open-source simulators for closed-loop benchmarks.
@@ 3.1.1 Parameter Initialization
Simulation offers the benefit of a high degree of control over the environment, including weather and lighting conditions, maps and 3D assets, and low-level attributes such as the arrangement and pose of objects in a traffic scene. While powerful, the number of these parameters is substantial, resulting in a challenging design problem. Current simulators tackle this problem in two ways:
Procedural Generation: Traditionally, initial parameters are hand-tuned by 3D artists and engineers, which we refer to as procedural generation [88, 89, 77, 91]. Each property is typically sampled from a probabilistic distribution with manually set parameters, which is a time-consuming process requiring a lot of expertise. This limits scalability. Nevertheless, this remains one of the most commonly used methods of initialization. Procedural generation algorithms combine rules, heuristics, and randomization to create diverse road networks, traffic patterns, lighting conditions, and object placements [93, 94].
Data-Driven: Data-driven approaches for simulation initialization aim to learn the required parameters. Arguably the simplest data-driven approach to initialization is to sample directly from a log of real-world driving data [92, 14]. In this method, parameters such as road maps or traffic patterns are directly extracted from pre-recorded datasets. The advantage of log sampling is its ability to capture the natural variability present in real-world data, leading to more realistic simulation scenarios than procedural generation. However, it may not encompass rare or extreme situations that are critical for testing the robustness of the autonomous driving system. The initial parameters can be optimized to increase the representation of such scenarios [7, 34, 35]. Another advanced data-driven approach to initialization is generative modeling, where machine learning algorithms are utilized to learn the underlying structure and distributions of real-world data. These algorithms can then generate novel scenarios that resemble the real world but were not included in the original data [95, 96, 97, 98].
@@ 3.1.2 Traffic Simulation
Traffic simulation involves generating and positioning virtual entities within the environment with realistic motion [95, 99]. These entities often include vehicles (such as trucks, cars, motorcycles, bicycles, etc.) and pedestrians. Traffic simulators must account for the effects of speed, acceleration, braking, obstructions, and the behavior of other entities. Moreover, traffic light states must be periodically updated to simulate realistic city driving. There are two popular approaches for traffic simulation, which we describe below.
Rule-Based: Rule-based traffic simulators use pre-defined rules to generate the motion of the traffic entities [77]. This approach is straightforward to implement, but the resulting motion may not be highly realistic. The most prominent implementation of this concept is the Intelligent Driver Model (IDM) [100]. IDM is a car-following model that computes acceleration for each vehicle based on its current speed, the speed of the leading vehicle, and a desired safety distance. Although widely used, IDM may be inadequate to capture complex interactions in urban environments.
@@ 3.1.3 Sensor Simulation
Sensor simulation is crucial for evaluating an end-to-end self-driving system. This involves generating simulated raw sensor data, such as camera images or LiDAR scans that the driving system would receive from the environment from different viewpoints in the simulator [104, 105, 106, 107]. This process needs to take into account noise and occlusions in order to realistically assess the autonomous system’s performance. There are two branches of ideas in the literature concerning sensor simulation, as described below.
Graphics-Based: Recent computer graphics simulators use 3D models of the environment, along with vehicle and traffic entity models, to generate sensor data via approximations of physical rendering processes in the sensors [89, 77, 91]. For example, this can take into account occlusions, shadows, and reflections present in real-world environments while simulating camera images. However, the realism of graphics-based simulation is often subpar or comes at the cost of extremely heavy computation, making parallelization non-trivial [108]. It is closely tied to the quality of the 3D models as well as the approximations used in modeling the sensors. A comprehensive survey of graphics-based rendering for driving data is provided in [109].
Data-Driven: Data-driven sensor simulation leverages and adapts real-world sensor data to create a new simulation where both the ego vehicle and background traffic may move differently from the way they did in the original data [110, 111, 112]. One popular approach is the use of Neural Radiance Fields (NeRF) [113], which can generate novel views of a scene by learning an implicit representation of the scene’s geometry and appearance. These methods can produce more realistic sensor data than graphics-based approaches, but they have limitations such as high rendering times or requiring independent training for each scene being reconstructed [114, 115, 116, 107]. Another approach to data-driven sensor simulation is domain adaptation, which aims to minimize the distribution shift between real and graphics-based simulated sensor data [117]. Machine learning techniques, such as GANs or style transfer, can be employed to improve realism. Further details regarding these ideas can be found in Sec. 4.8.3.
@@ 3.1.4 Benchmarks
We provide a succinct overview of driving benchmarks available up to date in Table I. In 2019, the original benchmark released with CARLA [77] was solved with near-perfect scores [5]. The subsequent NoCrash benchmark [118] involves training on a single CARLA town (Town01) under specific weather conditions and testing generalization to another town and set of weathers. Instead of a single town, the Town05 benchmark [6] involves training on all available towns in CARLA while withholding Town05 for testing. Similarly, the LAV benchmark trains on all towns except for Town02 and Town05, which are both reserved for testing to increase the diversity of the test routes. Roach [21] expands to a setting with 3 test towns, albeit all seen during training, and without the safety-critical scenarios used in Town05 and LAV. Finally, the Longest6 benchmark [28] uses 6 test towns. Two online submission servers for CARLA agents, known as the leaderboard (v1 and v2), are available at [13]. The leaderboard ensures fair comparisons by keeping the evaluation routes confidential. Leaderboard v2 is highly challenging due to the extremely long route length (over 8km on average, as opposed to 1-2km on v1) and a wide variety of new traffic scenarios. No methods have been benchmarked there yet.
@@
The nuPlan simulator is not currently accessible for end-to-end systems since the sensor data and corresponding sensor simulation aspect (Sec. 3.1.3) have not yet been made public. However, there are two existing benchmarks on which agents directly input maps and object properties available via the data-driven parameter initialization for nuPlan (Sec. 3.1.1). Val14, proposed in [119], uses a publicly accessible validation split of nuPlan [14]. The leaderboard, a submission server for testing on the private test set, was used in the 2023 nuPlan challenge. Unfortunately, this is no longer publicly available for submissions.
3.2 Offline Evaluation (Open-loop)
@@
Open-loop evaluation involves assessing a system’s performance against pre-recorded expert driving behavior. This method requires evaluation datasets that include (1) sensor readings, (2) goal locations, and (3) corresponding future driving trajectories, usually obtained from a human driver. Given sensor inputs and goal locations from the dataset as inputs, performance is measured by comparing the system’s predicted future trajectory against the trajectory taken by the human in the driving log. Systems are evaluated based on how closely their trajectory predictions match the human ground truth, as well as auxiliary metrics such as the collision probability with other agents [120]. The advantage of open-loop evaluation is that it is easy to implement and realistic traffic and sensor data, as it does not require a simulator. However, the key disadvantage is that it does not measure performance in the actual test distribution encountered by the system during deployment. During testing, the driving system may deviate from the expert driving corridor, and it is essential to verify the system’s ability to recover from such drift. Furthermore, the distance between the predicted trajectory and the observed trajectory is not a suitable indicator of performance in a multi-modal scenario. For example, in the case of merging into a turning lane, both the options of merging immediately or later are equally valid, but open-loop evaluation penalizes the option that was not observed in the data. Similarly, the predicted trajectory could depend on observations that will only be available in the future, e.g., stopping at a light that is still green but will turn red soon, a situation that is invalid to evaluate with a single ground truth trajectory.
@@
This approach requires a comprehensive dataset of trajectories to draw from. The most popular datasets for this purpose include nuScenes [121], Argoverse [122, 123], Waymo [124], and nuPlan [14]. All of these datasets comprise a large number of annotated trajectories from real-world driving environments with varying degrees of difficulty. However, open-loop results do not provide conclusive evidence of improved driving behavior in closed-loop, due to the aforementioned drawbacks [118, 120, 119]. Overall, a realistic closed-loop benchmarking, if available and applicable, is recommended in future research.
@@ 4 Challenges
For each topic/problem illustrated in Fig. 1, we now discuss the related works, current challenges, and promising future trends and opportunities. We start with challenges related to handling different input modalities and formulations in Sec. 4.1, followed by a discussion on visual abstraction for efficient policy learning in Sec. 4.2. Further, we introduce learning paradigms such as world model learning (Sec. 4.3), multi-task frameworks (Sec. 4.4), and policy distillation (Sec. 4.5). Finally, we discuss general issues that impede safe and reliable end-to-end autonomous driving, including interpretability in Sec. 4.6, causal confusion in Sec. 4.7, and robustness and generalizability in Sec. 4.8.
@@
Figure 4: Examples of input modality and fusion strategy. Different modalities have distinct characteristics, leading to the challenge of effectively fusing them and attending to action-critical features. We take point clouds and images as examples to depict various fusion strategies.
4.1 Input Modality
4.1.1 Multi-sensor Fusion
@@
Though early works [3, 8] successfully achieved simple self-driving tasks such as lane-following with a monocular camera, this single input modality is insufficient for handling complex scenarios [28]. Therefore, various sensors have been introduced and equipped on recent self-driving vehicles, as shown in Fig. 4. Particularly, RGB images from cameras naturally replicate how humans perceive the world with abundant semantic visual information; LiDARs or stereo cameras provide accurate 3D spatial knowledge. Additionally, vehicle states such as speed and acceleration from speedometers and IMUs, together with high-level navigation commands, are other lines of input that guide the end-to-end system. However, various sensors have distinct perspectives and data distributions, and the large gaps between them present great challenges in effectively fusing them to complement each other for autonomous driving.
@@
Multi-sensor fusion has predominantly been discussed in perception-related fields, e.g., object detection [125, 126], tracking [127, 128], and semantic segmentation [129, 130], and is typically categorized into three groups: early, mid, and late fusion. End-to-end autonomous driving algorithms explore similar fusion schemes [131]. Early fusion means combining sensory information before feeding it into the feature extractor. Concatenation is a common way for fusing various inputs, such as images and depth [132], BEV point clouds and HD maps [32, 133], etc., and then processing them with a shared feature extractor. [134, 135] paint LiDAR points on BEV with the same size as perspective-view images and combine them as inputs. To resolve the view discrepancy, some works attempt to project point clouds on 2D images [136] or attach an additional channel for each LiDAR point by predicting semantic labels from images in advance [137, 52]. On the other hand, the late fusion scheme combines multiple results from multi-modalities. It is less discussed due to its inferior performance [131, 6].
Contrary to these approaches, middle fusion achieves multi-sensor fusion within the network by separately encoding inputs and then combining them at the feature level. Naive concatenation is also frequently adopted to fuse features from different modalities [138, 139, 15, 140, 141, 142, 22, 30]. Recently, works have employed Transformers [27] to model interactions between pairs of features. Transfuser [6, 28] processes images and LiDAR inputs using two independent convolutional encoders, interconnecting each resolution of features with Transformer encoders, resulting in four-stage feature fusion. Self-attention layers are employed for sensor tokens, attending regions of interest and updating information from other modalities. MMFN [143] further incorporates OpenDrive map and radar inputs on top of Transfurser. [29, 144] take a one-stage Transformer encoder architecture to fuse various features after the last encoder block. The attention mechanism has demonstrated great effectiveness in aggregating the context of different sensor inputs and achieving safer end-to-end driving performance.
@@
Different modalities generally bring increased fields of view and perceptual accuracy, but fusing them to extract critical information for end-to-end autonomous driving requires further exploration. It is essential to model these modalities in a unified space such as BEV [125, 126], identify policy-related context [6, 12], and discard irrelevant sensory information. Besides, fully utilizing the powerful Transformer architecture remains a challenge. The self-attention layer interconnects all tokens to model their areas of interest freely, but it incurs a significant computational cost and cannot guarantee useful information extraction. More advanced Transformer-based multi-sensor fusion mechanisms in the perception field, such as [145, 146], hold promise for application to the end-to-end driving task.

### v2

@@ 3 Benchmarking
Autonomous driving systems require a comprehensive evaluation to ensure safety. Researchers must benchmark these systems using appropriate datasets, simulators, metrics, and hardware to accomplish this. This section delineates three approaches for benchmarking end-to-end autonomous driving systems: (1) real-world evaluation, (2) online or closed-loop evaluation in simulation, and (3) offline or open-loop evaluation on driving datasets. We focus on the scalable and principled online simulation setting and summarize real-world and offline assessments for completeness.
3.1 Real-world Evaluation
Early efforts on benchmarking self-driving involved real-world evaluation. Notably, DARPA initiated a series of races to advance autonomous driving. The first event offered \$1M in prize money for autonomously navigating a 240km route through the Mojave desert, which no team achieved [85]. The final series event, called the DARPA Urban Challenge, required vehicles to navigate a 96km mock-up town course, adhering to traffic laws and avoiding obstacles [86]. These races fostered important developments in autonomous driving, such as LiDAR sensors. Following this spirit, the University of Michigan established MCity [87], a large controlled real-world environment designed to facilitate testing autonomous vehicles. Besides such academic ventures, industries with the resources to deploy fleets of driverless vehicles also rely on real-world evaluation to benchmark improvements in their algorithms.
3.2 Online/Closed-loop Simulation
Conducting tests of self-driving systems in the real world is costly and risky. To address this challenge, simulation is a viable alternative [88, 89, 90, 91, 92, 14]. Simulators facilitate rapid prototyping and testing, enable the quick iteration of ideas, and provide low-cost access to diverse scenarios for unit testing. In addition, simulators offer tools for measuring performance accurately. However, their primary disadvantage is that the results obtained in a simulated environment do not necessarily generalize to the real world (Sec. 4.9.3).
Closed-loop evaluation involves building a simulated environment that closely mimics a real-world driving environment. The evaluation entails deploying the driving system in simulation and measuring its performance. The system has to navigate safely through traffic while progressing toward a designated goal location. There are four main sub-tasks involved in developing such simulators: parameter initialization, traffic simulation, sensor simulation, and vehicle dynamics simulation. We briefly describe these sub-tasks below, followed by a summary of currently available open-source simulators for closed-loop benchmarks.
@@ 3.2.1 Parameter Initialization
Simulation offers the benefit of a high degree of control over the environment, including weather, maps, 3D assets, and low-level attributes such as the arrangement of objects in a traffic scene. While powerful, the number of these parameters is substantial, resulting in a challenging design problem. Current simulators tackle this in two ways:
Procedural Generation: Traditionally, initial parameters are hand-tuned by 3D artists and engineers [88, 89, 90, 91]. This limits scalability. Recently, some of the simulation properties can be sampled from a probabilistic distribution with computer algorithms, which we refer to as procedural generation [93]. Procedural generation algorithms combine rules, heuristics, and randomization to create diverse road networks, traffic patterns, lighting conditions, and object placements [94, 95]. Due to its efficiency compared to fully manual design, it has become one of the most commonly used methods of initialization for video games and simulations. Nevertheless, the process still needs pre-defined parameters and algorithms to control generation reliability, which is time-consuming and requires a lot of expertise.
Data-Driven: Data-driven approaches for simulation initialization aim to learn the required parameters. Arguably, the simplest way is to sample from real-world driving logs [92, 14], where parameters such as road maps or traffic patterns are directly extracted from pre-recorded datasets. The advantage of log sampling is its ability to capture the natural variability present in real-world data, leading to more realistic simulation scenarios. However, it may not encompass rare situations that are critical for testing the robustness of autonomous driving systems. The initial parameters can be optimized to increase the representation of such scenarios [7, 34, 35]. Another advanced data-driven approach to initialization is generative modeling, where machine learning algorithms are utilized to learn the underlying structure and distributions of real-world data. They can then generate novel scenarios that resemble the real world but were not included in the original data [96, 97, 98, 99].
@@ 3.2.2 Traffic Simulation
Traffic simulation involves generating and positioning virtual entities in the environment with realistic motion [97, 100]. These entities often include vehicles (such as cars, motorcycles, bicycles, etc.) and pedestrians. Traffic simulators must account for the effects of speed, acceleration, braking, obstructions, and the behavior of other entities. Moreover, traffic light states must be periodically updated to simulate realistic city driving. There are two popular approaches for traffic simulation, which we describe below.
Rule-Based: Rule-based traffic simulators use pre-defined rules to generate the motion of traffic entities. The most prominent implementation of this concept is the Intelligent Driver Model (IDM) [101]. IDM is a car-following model that computes acceleration for each vehicle based on its current speed, the speed of the leading vehicle, and a desired safety distance. Although widely used and straightforward, this approach may be inadequate to simulate realistic motion and complex interactions in urban environments.
@@ 3.2.3 Sensor Simulation
Sensor simulation is crucial for evaluating end-to-end self-driving systems. This involves generating simulated raw sensor data, such as camera images or LiDAR scans that the driving system would receive from different viewpoints in the simulator [105, 106, 107]. This process needs to take into account noise and occlusions to realistically assess the autonomous system. There are two main branches of ideas concerning sensor simulation, as described below.
Graphics-Based: Recent computer graphics simulators use 3D models of the environment, along with traffic entity models, to generate sensor data via approximations of physical rendering processes in the sensors [89, 90]. For example, this can involve occlusions, shadows, and reflections present in real-world environments while simulating camera images. However, the realism of graphics-based simulation is often subpar or comes at the cost of heavy computation, making parallelization non-trivial [108]. It is closely tied to the quality of the 3D models and the approximations used in modeling the sensors. A comprehensive survey of graphics-based rendering for driving data is provided in [109].
Data-Driven: Data-driven sensor simulation leverages real-world sensor data to create the simulation where both the ego vehicle and background traffic may move differently from the way they did in recordings [110, 111, 112]. Popular methods are Neural Radiance Fields (NeRF) [113] and 3D Gaussian Splatting [114], which can generate novel views of a scene by learning an implicit representation of the scene’s geometry and appearance. These methods can produce more realistic sensor data visually than graphics-based approaches, but they have limitations such as high rendering times or requiring independent training for each scene being reconstructed [115, 116, 117, 118, 107]. Another approach to data-driven sensor simulation is domain adaptation, which aims to minimize the gap between real and graphics-based simulated sensor data [119]. Deep learning techniques such as GANs can be employed to improve realism (Sec. 4.9.3).
3.2.4 Vehicle Dynamics Simulation
The final aspect of driving simulation pertains to ensuring that the simulated vehicle adheres to physically plausible motion. Most existing publicly available simulators use highly simplified vehicle models, such as the unicycle model [120] or the bicycle model [121]. However, in order to facilitate seamless transfer of algorithms from simulation to the real world, it is essential to incorporate more accurate physical modeling of vehicle dynamics. For instance, CARLA adopts a multi-body system approach, representing a vehicle as a collection of sprung masses on four wheels. For a comprehensive review, please refer to [122].
@@ 3.2.5 Benchmarks
We give a succinct overview of end-to-end driving benchmarks available up to date in Table I. In 2019, the original benchmark released with CARLA [90] was solved with near-perfect scores [5]. The subsequent NoCrash benchmark [123] involves training on a single CARLA town under specific weather conditions and testing generalization to another town and set of weathers. Instead of a single town, the Town05 benchmark [6] involves training on all available towns while withholding Town05 for testing. Similarly, the LAV benchmark trains on all towns except Town02 and Town05, which are both reserved for testing. Roach [21] uses a setting with 3 test towns, albeit all seen during training, and without the safety-critical scenarios in Town05 and LAV. Finally, the Longest6 benchmark [28] uses 6 test towns. Two online servers, the leaderboard (v1 and v2) [13], ensure fair comparisons by keeping evaluation routes confidential. Leaderboard v2 is highly challenging due to the long route length (over 8km on average, as opposed to 1-2km on v1) and a wide variety of new traffic scenarios.
@@
The nuPlan simulator is currently accessible for evaluating end-to-end systems via the NAVSIM project [124]. Further, there are two benchmarks on which agents input maps and object properties via the data-driven parameter initialization for nuPlan (Sec. 3.2.1). Val14, proposed in [125], uses a validation split of nuPlan. The leaderboard, a submission server with the private test set, was used in the 2023 nuPlan challenge, but it is no longer public for submissions.
3.3 Offline/Open-loop Evaluation
@@
Open-loop evaluation mainly assesses a system’s performance against pre-recorded expert driving behavior. This method requires evaluation datasets that include (1) sensor readings, (2) goal locations, and (3) corresponding future driving trajectories, usually obtained from human drivers. Given sensor inputs and goal locations as inputs, performance is measured by comparing the system’s predicted future trajectory against the trajectory in the driving log. Systems are evaluated based on how closely their trajectory predictions match the human ground truth, as well as auxiliary metrics such as the collision probability with other agents. The advantage of open-loop evaluation is that it is easy to implement using realistic traffic and sensor data, as it does not require a simulator. However, the key disadvantage is that it does not measure performance in the actual test distribution encountered during deployment. During testing, the driving system may deviate from the expert driving corridor, and it is essential to verify the system’s ability to recover from such drift (Sec. 4.9.2). Furthermore, the distance between the predicted and the recorded trajectories is not an ideal metric in a multi-modal scenario. For example, in the case of merging into a turning lane, both the options of merging immediately or later could be valid, but open-loop evaluation penalizes the option that was not observed in the data. Therefore, besides measuring collision probability and prediction errors, a few metrics were proposed to cover more comprehensive aspects such as traffic violations, progress, and driving comfort [125].
@@
This approach requires comprehensive datasets of trajectories to draw from. The most popular datasets for this purpose include nuScenes [126], Argoverse [127], Waymo [128], and nuPlan [14]. All of these datasets comprise a large number of real-world driving traversals with varying degrees of difficulty. However, open-loop results do not provide conclusive evidence of improved driving behavior in closed-loop, due to the aforementioned drawbacks [123, 129, 125, 130]. Overall, a realistic closed-loop benchmarking, if available and applicable, is recommended in future research.
@@ 4 Challenges
Following each topic illustrated in Fig. 1, we now walk through current challenges, related works or potential resolutions, risks, and opportunities. We start with challenges in handling different input modalities in Sec. 4.1, followed by a discussion on visual abstraction for efficient policy learning in Sec. 4.2. Further, we introduce learning paradigms such as world model learning (Sec. 4.3), multi-task frameworks (Sec. 4.4), and policy distillation (Sec. 4.5). Finally, we discuss general issues that impede safe and reliable end-to-end autonomous driving, including interpretability in Sec. 4.6, safety guarantees in Sec. 4.7, causal confusion in Sec. 4.8, and robustness in Sec. 4.9.
@@
Figure 4: Examples of input modality and fusion strategy. Different modalities have distinct characteristics, leading to the challenge of effective sensor fusion. We take point clouds and images as examples to depict various fusion strategies.
4.1 Dilemma over Sensing and Input Modalities
4.1.1 Sensing and Multi-sensor Fusion
@@
Sensing: Though early work [8] successfully achieved following a lane with a monocular camera, this single input modality cannot handle complex scenarios. Therefore, various sensors in Fig. 4 have been introduced for recent self-driving vehicles. Particularly, RGB images from cameras replicate how humans perceive the world, with abundant semantic details; LiDARs or stereo cameras provide accurate 3D spatial knowledge. Emerging sensors like mmWave radars and event cameras excel at capturing objects’ relative movement. Additionally, vehicle states from speedometers and IMUs, together with navigation commands, are other lines of input that guide the driving system. However, various sensors possess distinct perspectives, data distributions, and huge price gaps, thereby posing challenges in effectively designing the sensory layout and fusing them to complement each other for autonomous driving.
@@
Multi-sensor fusion has predominantly been discussed in perception-related fields, e.g., object detection [131, 132] and semantic segmentation [133, 134], and is typically categorized into three groups: early, mid, and late fusion. End-to-end autonomous driving algorithms explore similar fusion schemes. Early fusion combines sensory inputs before feeding them into shared feature extractors, where concatenation is a common way for fusion [135, 136, 137, 32, 138]. To resolve the view discrepancy, some works project point clouds on images [139] or vice versa (predicting semantic labels for LiDAR points [140, 52]). On the other hand, late fusion combines multiple results from multi-modalities. It is less discussed due to its inferior performance [141, 6]. Contrary to these methods, middle fusion achieves multi-sensor fusion within the network by separately encoding inputs and then fusing them at the feature level. Naive concatenation is also frequently adopted [142, 143, 15, 144, 145, 146, 22, 30]. Recently, works have employed Transformers [27] to model interactions among features  [6, 28, 147, 29, 148]. The attention mechanism in Transformers has demonstrated great effectiveness in aggregating the context of different sensor inputs and achieving safer end-to-end driving.
@@
Inspired by the progress in perception, it is beneficial to model modalities in a unified space such as BEV [131, 132]. End-to-end driving also requires identifying policy-related contexts and discarding irrelevant details. We discuss perception-based representations in Sec. 4.2.1. Besides, the self-attention layer, interconnecting all tokens freely, incurs a significant computational cost and cannot guarantee useful information extraction. Advanced Transformer-based fusion mechanisms in the perception field, such as [149, 150], hold promise for application to the end-to-end driving task.